OpenStudy (anonymous):

A simple random sample of voters is taken from the voters in a large state. Using the methods of our course, researchers construct an approximate 99% confidence interval for the percent of the state’s voters who will vote for Candidate A. The interval goes from 37.3% to 48.7%. In the sample, the percent of voters who will vote for Candidate A is equal to ____________%.

OpenStudy (amistre64):

well, since an interval is centered around the sample mean .... what is the midpoint between 37.3 and 48.7?

OpenStudy (amistre64):

even if they are using "approximation" ; that is still 3 standard deviations centered about the sample mean

OpenStudy (anonymous):

The middle point is (37.3+48.7)/2 =43

OpenStudy (anonymous):

For A, I would find the center of 37.3% to 48.7%. i.e. (37.3 + 48.7)/2 = 43.0 the interval from 43 to 48.7 = 5.7 represents 2.58 σ so 5.7/2.58 = 2.21 % per σ

OpenStudy (anonymous):

is correct the statement?

OpenStudy (amistre64):

\[p\pm E\] \[37.3~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~48.7\\~~~|<------p------>|\\~~~~~~~~~~~~~-E~~~~~~~~~~~~~~~~~~~~+E\] the sample percentage is the midpoint between the extremes

OpenStudy (amistre64):

add the ends and divide in half \[\frac{37.3+48.7}{2}=\frac{86.0}{2}=43\]

OpenStudy (amistre64):

43% of the sample choose A

OpenStudy (amistre64):

finding the sample size would be fun :)

OpenStudy (anonymous):

Exercise set 5 answers: http://4c51a450.linkbucks.com